# Application of Battery Storage to Switching Predictive Control of Power Distribution Systems Including Road Heating*

### v1

@@ Abstract
A road heating system is an electrical device which promotes snow melting by burying a heating cable as a thermal source underground. When integrating road heating into the power distribution system, we need to optimize the flow of electric power by appropriately integrating distributed power sources and conventional power distribution equipment. In this paper, we extend the power distribution system considered in the authors’ previous study to the case where battery storage is installed. As a main result, we propose a predictive switching control that achieves the reduction of distribution loss, attenuation of voltage fluctuation, and efficient snow melting, simultaneously. We verify the effectiveness of the application of battery storage through numerical simulation.
@@ I Introduction
In heavy snowfall regions, there has long been a serious problem of great inconvenience and danger to people’s lives, for example falling on sidewalks, heavy workload due to snow removal, traffic congestion and etc [1][2]. A road heating system is one of the electrical devices which promotes snow melting by burying a thermal source underground, and raises the temperature of the ground surface [3]. However, the road heating system consumes a large amount of electrical power when we integrate road heating into the power distribution system [4][5]. We need to optimize the flow of electrical power by properly integrating distributed power sources, e.g. photovoltaic power generation (PV), and electric vehicles (EVs) with conventional power distribution equipment for an appropriate operation of power distribution systems.
@@
Against the problem explained in the above paragraph, the authors of this paper [6] have derived a mathematical model of the power distribution system including road heating, and of snowmelting on the ground surface via thermal diffusion in the underground based on the nonlinear ordinary differential equation (ODE) model [7][8]. Furthermore, the authors proposed a predictive control by switching of the power distribution system including road heating [9]. The proposed control of  [9] achieves reduction of power distribution losses, attenuation of voltage impact, and efficient snowmelting, simultaneously. However, the supply of power electricity from the PV power generation equipment is limited to the distribution system only. There is a potential for improvement in terms of power flow optimization.
@@
Based on the considerations given in the above paragraph, the purpose of this paper is to extend the results of the previous study [9] in the following two points.
  • We introduce a switching to the PV power generation equipment to achieve a more efficient power distribution system operation which does not depend on transmission from the grid.
@@
  • A battery storage is introduced to the considered power distribution system to make effective use of surplus power generated by PV in the past.
Based on the above extension, this paper improves the optimization of power flow by applying the battery storage to the power distribution system, making more efficient use of the available power from the PV power generation equipment and the battery storage. As a main result, we give a design of predictive switching control of the distribution system under study, as well as a design of predictive control of the PV power generation equipment and battery storage. Finally, we verify the effectiveness of these predictive controls through numerical simulations. Through this verification, we show that the predictive control of the distribution system is more efficient than the previous study  [9].
@@
The outline of this paper is described as follows. In Section II, we introduce a mathematical model of a power distribution system including road heating. As a main result, we give a switching predictive control for the system based on an evaluation function to simultaneously achieve reduction of distribution losses, attenuation of voltage impact, and efficient snowmelting in Section III. Finally, we verify an effectiveness of the proposed switching predictive control via numerical simulation in Section IV.
@@ II Modeling of Power Distribution System
In this section, we develop the distribution system including road heating considered in a previous study due to the authors [6][9] to a mathematical model for the case including switching of the discharging of PV power generation equipment and battery storage.
@@ II-A Overall setup of system
In this subsection, we describe the overall setup of the power distribution system including road heating. We show an overview of the system in Fig. 1. In the system, electrical power is supplied from the underground power distribution lines to the load electrical equipment (residences, EVs, lamps, and etc.) and from the PV electric power generation equipment to the underground power distribution lines. Assuming a realistic burial location, a heating cable is buried in parallel with the underground power distribution line in the horizontal direction. Two switches are connected at both terminals to control the voltage distribution over the heating cable. When Switch 1 is turned On, the power electricity is transmitted from the starting end of the heating cable. On the other hand, the power is transmitted from the battery storage if Switch 2 is turned On. In addition, the heat generated in the cable of the distribution line is transferred within the cable, and diffused into the underground. The Joule heat propagating to the ground surface is used to melt snow.
@@
We show a single-line diagram of the power distribution system in Fig. 2. The branches of the underground power distribution line and the heating cable at the starting point and terminal are denoted as Bifurcation points 1 and 2, respectively. We assume that Transformers A and B are connected to Switches 1 and 2, respectively.
@@ II-B Voltage distribution on underground power distribution line [7][8]
II-B1 Voltage distribution on underground power distribution line
We describe the voltage distribution of an underground power distribution line, which is assumed to be in three-phase equilibrium, by the ordinary differential equations with position $x$ [m] as an independent variable, for an arbitrarily fixed time $t$ [min]. We denote the voltage phasor for single-phase at position $x$ and time $t$ by $v_{\rm e}(x;t)e^{j\theta_{\rm e}(x;t)}$, where $j$ is the imaginary unit. Then, $\theta_{\rm e}(x;t)$ [rad] and $v_{\rm e}(x;t)$ [V] correspond to the voltage phase and amplitude, respectively. The spatial variation on the distribution voltage profile is described by the nonlinear ODE model [7][8]:
@@ II-B1 Voltage distribution on underground power distribution line
where $s_{\rm e}(x,t)$ [V2/m] and $w_{\rm e}(x,t)$ [V/m] represent the supplemental variable and voltage gradient at position $x$ and time $t$, respectively. In (1) and (2), $p_{\rm e}(x,t)$ [W/m] and $q_{\rm e}(x,t)$ [Var/m] are active and reactive powers at position $x$ and time $t$ of the underground power distribution line, respectively. If $p_{\rm e}(x,t)>0$ holds, $p_{\rm e}(x,t)$ denotes the active power, i.e. the supply, flowing to the distribution line at $x$. On the other hand, if $p_{\rm e}(x,t)<0$ holds, $p_{\rm e}(x,t)$ corresponds to the active power consumption flowing from the line at $x$. The same inequality holds for the reactive power $q_{\rm e}(x,t)$. Due to the assumptions made at the terminal of the underground power distribution line, $p_{\rm e}(L_{\rm f},t)=0$ and $q_{\rm e}(L_{\rm f},t)=0$ hold. The constants $g_{\rm e}$ [S/m] and $b_{\rm e}$ [S/m] are the conductance and susceptance per unit length at position $x$ of the line. Since the distribution loss $\Gamma_{\rm e}(x;t)$ [W/m] of the underground distribution line is considered for three phases, the distribution loss for a single-phase is three times larger, and is given by
from [12].
@@ II-B2 Boundary condition
The voltage phase and amplitude satisfy the following boundary conditions at each point:
@@
In (3), $\theta_{1}$ and $v_{1}$ denote the voltage phase and amplitude in Bifurcation point 1, respectively, which are supposed to be constant. Moreover, $\theta_{2}(t)$ and $v_{2}(t)$ are the voltage phase and amplitude in Bifurcation point 2 at time $t$, respectively, in (4).
@@ II-C Voltage and temperature distribution of heating cable [6]
We assume that the heating cable is single-phase two-wire system. The heating cable extends in a straight line from Transformer A to Transformer B with the switches connected at both ends. The horizontal spatial variable of the heating cable is $x$ [m] in common with underground power distribution line. The vertical spatial variable $y$ [cm] is the depth from the ground surface to the heating cable. This implies that $y=0$ [cm] and $y=D_{\rm f}$ [cm] correspond to the ground and the surface, respectively.
@@ II-C1 Voltage distribution
We denote the voltage phasor for single-phase at position $x$ and time $t$ by $v_{\rm h}(x;t)e^{j\theta_{\rm h}(x;t)}$, where $\theta_{\rm h}(x;t)$ [rad] and $v_{\rm h}(x;t)$ [V] are the voltage phase and amplitude, respectively. Similarly to the underground power distribution line, the spatial variation of the voltage distribution over the heating cable, which is assumed to be single-phase two-wire system, is described by the nonlinear ODE model [6].
@@
II-C2 Setting of transformers
We define $\theta_{\circ}(t)$ [rad], $v_{\circ}(t)$ [V], $s_{\circ}(t)$ [V2], and $w_{\circ}(t)$ [V/m] as the voltage phase, voltage amplitude, supplemental variable, and voltage gradient of Transformer $\circ$ $(\circ=\mathrm{A},\mathrm{B})$. At Transformer A, $\theta_{\mathrm{A}}(t)=\theta_{\mathrm{A}}$ and $v_{\mathrm{A}}(t)=v_{\mathrm{A}}$ are supposed to be constant. The turn ratio corresponds to the transformation ratio of the voltage amplitudes. The following equalities hold at each transformer from [8]:
Transformer A: $\theta_{\mathrm{h}}(0,t)=\theta_{\mathrm{A}},\ v_{\mathrm{h}}(0,t)=\frac{1}{a_{\rm A}}v_{\mathrm{A}},$
$s_{\mathrm{h}}(0,t)=s_{\mathrm{A}}(t),\ w_{\mathrm{h}}(0,t)=a_{\rm A}w_{\mathrm{A}}(t).,$
Transformer B: $\theta_{\mathrm{h}}(L_{\mathrm{f}},t)=\theta_{\mathrm{B}}(t),\ v_{\mathrm{h}}(L_{\mathrm{f}},t)=\frac{1}{a_{\mathrm{B}}}v_{\mathrm{B}}(t),$
$s_{\mathrm{h}}(L_{\mathrm{f}},t)=s_{\mathrm{B}}(t),\ w_{\mathrm{h}}(L_{\mathrm{f}},t)=a_{\rm B}w_{\mathrm{B}}(t).$
@@ II-C2 Setting of transformers
In the above conditions, $\theta_{1}$ [rad] and $v_{1}$ [V] denote the reference voltage phase and amplitude for the underground distribution line at $x=0$. Moreover, $a_{\rm 1}$ is the transformer ratio of the single phase transformer.
@@ II-C3 Thermal distribution of heating cable surface
$\frac{\mathrm{d}\delta_{\rm surf,diff}(t;x)}{\mathrm{d}t}=\frac{\Gamma_{\rm h}(x,t)\times 10^{-2}-q_{\rm r}-\gamma_{\rm cable}\delta_{\rm surf,diff}(t;x)}{C_{\rm cable}}$
@@ II-D Switches
Since the voltages at Transformer B and the terminal $(x=L_{\rm f})$ of the underground power distribution line are required to be equal, if one switch is On, the other switch is Off. Moreover, if the heating cable is not used, both switches are Off. We define $\sigma_{i}(t)\in\{0,1\}$ $(i=0,1,2)$ as the binary variable by
@@
$\sigma_{{\rm pv}}(t)$ :=$:=\begin{cases}1&\mbox{(Underground distribution line)}\\ 0&\mbox{(Battery storage)}.\\ \end{cases}$
@@
$\sigma_{{\rm battery},1}(t)$ :=$:=\begin{cases}1&\mbox{(Underground distribution line)}\\ 0&\mbox{(Otherwise), }\\ \end{cases}$
@@ III Switching Predictive Control
In this section, we explain the predictive control by switches connected to the heating cables and the predictive control of the PV power generation equipment and the battery storage as the main results of this paper.
@@ III-A Design of Switching Predictive Control
In this susection, we explain the overall process for the predictive switching control of the power distribution system, the PV power generation equipment, and predictive control of battery storage. The optimal control inputs are computed according to Steps 0-5 below. Note that Step 0 and Steps 1-4 correspond to the preliminary and during operation of the predictive control, respectively. We also add Steps 3 and 4 to the predictive switching control of the previous study [9] due to the authors of this paper.
@@
  • Step 1: We define $N:=\frac{T_{\rm pred}}{T_{\rm mini}}$ as the total number of times at which switching can occur. This number corresponds to the sum of all $3^{N}$ switching patterns, $2^{N}$ PV transmission patterns, and $3^{N}$ discharging patterns of the battery storage. For all possible $3^{N}+2^{N}+3^{N}$ patterns, we perform the simulations from $t=kT_{\rm mini}$ to $t+T_{\rm pred}=(k+N)T_{\rm mini}$ based on the mathematical model of the distribution system.
  • Step 2: Based on the simulation in Step 1, we compute the value of the evaluation function $J[k]$ for the possible $3^{N}$ switching patterns ${\bf\it\mathbf{\sigma}}_{\rm d}[k]$ at $t=kT_{\rm mini}$. We then compare the $3^{N}$ values of the functions. We choose the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}[k]$ that has the smallest value at $t=kT_{\rm mini}$. We explain the detailed procedure of the design of the optimal switching pattern in Section III-B.
  • Step 3: Based on the simulation and the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}^{*}[k]$ computed in Steps 1 and 2, respectively, we compute the evaluation function for the $2^{N}$ possible transmission patterns of the PV power generation equipment at $t=kT_{\rm mini}$. We compute the value of the evaluation function $J_{\rm pv}[k]$ for ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ of the transmission patterns of the PV power generation equipment at $t=kT_{\rm mini}$. Then, by comparing the $2^{N}$ values of the evaluation functions, we find the optimal transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ of the PV power generation equipment whose value is the minimum at $t=kT_{\rm mini}$. The detailed procedure of this step is explained in Section III-C.
  • Step 4: Based on the predictive simulation in Step 1, the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}^{*}[k]$ in Step 2, and the optimal transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ in Step 3, we compute the values of the evaluation function $J_{\rm battery}[k]$ for the $3^{N}$ battery storage transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]$ at $t=kT_{\rm mini}$ for the evaluation function $J_{\rm battery}[k]$. In this case, the $3^{N}$ values of the evaluation function are compared, and the battery transfer pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]$ with the minimum value at $t=kT_{\rm mini}$ is selected as the optimal power transfer pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}^{*}[k]$. We explain the detailed procedure of the design of the optimal switching pattern in Section III-D.

### v2

@@ Abstract
In regions with heavy snowfall, the living environment is becoming a serious problem due to heavy snow accumulation. A road heating is an electrical device which promotes snow melting by burying a heating cable as a thermal source underground in such regions. When integrating the road heating into power distribution systems, we need to optimize the flow of electric power by appropriately integrating distributed power sources and conventional power distribution equipment. In this paper, we introduce a battery storage to the power distribution system including road heating, and extend the predictive switching control of the systems due to the authors’ previous study to the case where battery storage is installed. As a main result, we propose a predictive switching control that utilizes photovoltaic (PV) power generation and surplus power stored in the battery storage effectively, and achieves the reduction of distribution loss, attenuation of voltage fluctuation, and efficient snow melting, simultaneously. We verify the effectiveness of the application of battery storage through numerical simulation using actual time series data of weather conditions and active power of the PV power generation and load.
@@ I Introduction
In heavy snowfall regions, there has long been a serious problem of great inconvenience and danger to people’s lives, for example falling on sidewalks, heavy workload due to snow removal, traffic congestion and etc [1][2]. A road heating is one of the electrical devices which promotes snow melting by burying a thermal source underground, and raises the temperature of the ground surface [3]. However, the road heating consumes a large amount of electrical power when we integrate road heating into the power distribution system [4][5]. The flow of electrical power should be optimized by properly integrating distributed power sources, e.g. photovoltaic power generation (PV), and electric vehicles (EVs) with conventional power distribution equipment for an appropriate operation of power distribution systems.
@@
Against the problem in the above paragraph, the authors of this paper [6] have derived a mathematical model of the power distribution system including road heating, and of snow melting on the ground surface via thermal diffusion in the underground based on the nonlinear ordinary differential equation (ODE) model [7][8]. Based on this model, the authors proposed a switching predictive control of the power distribution system including road heating [9]. The proposed control of [9] achieves reduction of power distribution losses, attenuation of voltage fluctuation, and efficient snow melting, simultaneously. However, the supply of power electricity from the PV power generation equipment is limited to the distribution system only. There is a potential for improvement from a view point of power flow optimization. In [10], a 2-dimensional flat ground surface is considered as the target of snow melting, and snow melting is occured by heat generation from thermal pipes. This study is different from the framework of an electrical heating cable as part of an electric power distribution system as considered in this paper.
@@
In this paper, we extend the results of the previous study [9] in the following two points from the observations given in the above paragraph, .
@@
  • A battery storage is introduced to the considered power distribution system to make effective use of surplus power generated by the PV power generation in the past.
  • We introduce a switching to the PV power generation equipment for a more efficient operation of the power distribution system which does not depend on transmission from the grid.
Based on the above extension, we give a design of predictive switching control of the distribution system including the switching of the PV power generation equipment and battery storage as a main result. Note that the authors of this paper have reported preliminary result in the reference [11] toward the objective explained at the beginning of this paragraph. The results of this paper provide a mathematically more accurate description of the mathematical model of the considered power distribution system than the reference [11]. In addition, this paper provides an additional case on the weather conditions for the simulation verification, and give comparisons with the result for the case without the battery storage in the authors’ previous study [9].
@@
The outline of this paper is described as follows. In Section II, we introduce a mathematical model of the considered power distribution system including with switching of the PV power generation equipment and battery storage. As a main result, we give a switching predictive control for the system based on an evaluation function to simultaneously achieve reduction of distribution losses, attenuation of voltage impact, and efficient snow melting in Section III. Finally, we verify an effectiveness of the proposed switching predictive control and the application of the battery storage via numerical simulation using actual data of weather conditions and active power of PV and load in Section IV.
@@ II Modeling of Power Distribution System
In this section, we develop the distribution system including road heating considered in the previous studies due to the authors [6][9] to a mathematical model for the case including switching of the discharging of PV power generation equipment and battery storage. In Subsection II-A, we describe the overall power distribution system considered in this paper. We provide the mathematical models of the voltage and temperature distribution of the underground power distribution line and heating cable that constitute the distribution system in Subsections II-B and II-C, respectively. Furthermore, in Subsection II-D, we introduce the mathematical models of the branch points, switches and transformers that interconnect them. Finally, in Subsection II-E, we describe the mathematical models of thermal diffusion in the underground and the snow melting on the ground surface. The details of the mathematical model of the overall system is summarized in the authors’ previous studies [6][9].
@@ II-A Overall setup of system
In this subsection, we describe the overall setup of the power distribution system including road heating. We show an overview of the system in Fig. 1. In the system, electrical power is supplied from the underground power distribution line to the load electrical equipment (residences, EVs, lamps, and etc.) and from the PV power generation equipment to the underground power distribution line. Assuming a realistic burial location, a heating cable is buried in parallel with the underground power distribution line in the horizontal direction. Two switches are connected at both terminals to control the voltage distribution over the heating cable. When Switch 1 is turned On, the power electricity is transmitted from the starting end of the heating cable. On the other hand, the power is transmitted from the battery storage if Switch 2 is turned On. In addition, the heat generated in the cable of the distribution line is transferred within the cable, and diffused into the underground. The Joule heat propagating to the ground surface is used to melt snow. No load electrical equipment or PV power generation equipment is connected at the terminal of the underground power distribution line and the open ends of the heating cable. We suppose that the Joule heat generated by the underground power distribution line is sufficiently small compared to the heat due to the heating cable. We show a single-line diagram of the power distribution system in Fig. 2. The branches of the underground power distribution line and the heating cable at the starting point and terminal are denoted as Bifurcation points 1 and 2, respectively.
@@ II-B Voltage distribution on underground power distribution line [7][8]
II-B1 Voltage distribution
We describe the voltage distribution of the underground power distribution line, which is assumed to be in three-phase equilibrium, by the ODEs with position $x$ [m] as an independent variable, for an arbitrarily fixed time $t$ [min]. We denote the voltage phasor for single-phase at position $x$ and time $t$ by $v_{\rm e}(x;t)e^{j\theta_{\rm e}(x;t)}$, where $j$ is the imaginary unit. Then, $\theta_{\rm e}(x;t)$ [rad] and $v_{\rm e}(x;t)$ [V] correspond to the voltage phase and amplitude, respectively. The spatial variation on the distribution voltage profile is described by the nonlinear ODE model [7][8]:
@@ II-B1 Voltage distribution
where $s_{\rm e}(x,t)$ [V2/m] and $w_{\rm e}(x,t)$ [V/m] represent the supplemental variable and voltage gradient at position $x$ and time $t$, respectively. In (1) and (2), $p_{\rm e}(x,t)$ [W/m] and $q_{\rm e}(x,t)$ [Var/m] are active and reactive powers at position $x$ and time $t$ of the underground power distribution line, respectively. If $p_{\rm e}(x,t)>0$ holds, $p_{\rm e}(x,t)$ denotes the active power, i.e. the supply, flowing to the distribution line at $x$. On the other hand, if $p_{\rm e}(x,t)<0$ holds, $p_{\rm e}(x,t)$ corresponds to the active power consumption flowing from the line at $x$. The same inequality holds for the reactive power $q_{\rm e}(x,t)$. Due to the assumptions made at the terminal of the underground power distribution line, $p_{\rm e}(L_{\rm f},t)=0$ and $q_{\rm e}(L_{\rm f},t)=0$ hold. The constants $g_{\rm e}$ [S/m] and $b_{\rm e}$ [S/m] are the conductance and susceptance per unit length at position $x$ of the line.
@@ II-B2 Boundary condition
The voltage phase and amplitude of the underground power distribution line satisfy the following boundary conditions at each point:
@@
In (3), $\theta_{1}$ and $v_{1}$ denote the constant voltage phase and amplitude in Bifurcation point 1, respectively, which correspond to the reference voltage phase and amplitude for the line at $x=0$. Moreover, $\theta_{2}(t)$ and $v_{2}(t)$ are the voltage phase and amplitude in Bifurcation point 2 at time $t$, respectively, in (4).
@@ II-C Voltage and temperature distribution of heating cable [6]
We assume that the heating cable is single-phase two-wire system. The heating cable extends in a straight line from the transformer at the starting point, to the terminal, with the switches connected at both ends. The horizontal spatial variable of the heating cable is $x$ [m] in common with underground power distribution line. The vertical spatial variable $y$ [cm] is the depth from the ground surface to the heating cable. This implies that $y=0$ [cm] and $y=D_{\rm f}$ [cm] correspond to the ground and the surface, respectively. The length of the switches is negligible because it is supposed sufficiently short compared to the underground power distribution line and heating cable. No voltage regulation devices are installed on the intermediate positions over the cable similarly to the underground power distribution line.
@@ II-C1 Voltage distribution
We denote the voltage phasor for single-phase at position $x$ and time $t$ by $v_{\rm h}(x;t)e^{j\theta_{\rm h}(x;t)}$, where $\theta_{\rm h}(x;t)$ [rad] and $v_{\rm h}(x;t)$ [V] are the voltage phase and amplitude, respectively. Similarly to the underground power distribution line, the spatial variation of the voltage distribution over the heating cable, which is assumed to be single-phase two-wire system, is described by the nonlinear ODE model [6]:
@@
II-C2 Boundary condition
@@
In the above conditions, $a_{\rm 1}$ is the transformer ratio of the single phase transformer.
@@ II-C3 Thermal distribution of heating cable surface
$\frac{\mathrm{d}\delta_{\rm surf,diff}(t;x)}{\mathrm{d}t}$
$=\frac{\Gamma_{\rm h}(x,t)\times 10^{-2}-q_{\rm r}-\gamma_{\rm cable}\delta_{\rm surf,diff}(t;x)}{C_{\rm cable}}$
@@ II-D Switches
Since the voltages at Transformer B and the terminal $(x=L_{\rm f})$ of the underground power distribution line are required to be equal, if one switch is On, the other switch is Off. Moreover, if the heating cable is not used, both switches are Off. We define binary variable $\sigma_{i}(t)\in\{0,1\}$ $(i=0,1,2)$ by
@@
$\sigma_{{\rm pv}}(t)$ :=$:=\begin{cases}1&\mbox{(Underground power distribution line)}\\ 0&\mbox{(Battery storage)}.\\ \end{cases}$
@@
$\sigma_{{\rm battery},1}(t)$ :=$:=\begin{cases}1&\mbox{(Underground power distribution line)}\\ 0&\mbox{(Otherwise), }\\ \end{cases}$
@@ III Switching Predictive Control
In this section, we explain the predictive control by switches connected to the heating cables and the predictive control of the PV power generation equipment and the battery storage as the main result of this paper. At first, in Subsection III-A, we describe the overall design of the proposed predictive switching control. Next, in Subsection III-B, the evaluation function and optimal switching patterns are introduced for the switching of the heating cable. Moreover, a similar formulation is given for the PV power generation equipment in Subsection III-C. Finally, in Subsection III-D, a mathematical description is provided with respect to the optimal discharging of the battery storage.
@@ III-A Design of Switching Predictive Control
In this subsection, we explain the overall process for the predictive switching control of the power distribution system, the PV power generation equipment, and predictive control of battery storage. The optimal control inputs are computed according to Steps 0-5 below. Note that Step 0 and Steps 1-4 correspond to the preliminary and during operation of the predictive control, respectively. We also add Steps 3 and 4 to the predictive switching control of the previous study [9] due to the authors.
@@
  • Step 1: We define $N:=\frac{T_{\rm pred}}{T_{\rm mini}}$ as the total number of times at which switching can occur. This number corresponds to the sum of all $3^{N}$ switching patterns, $2^{N}$ PV transmission patterns, and $3^{N}$ discharging patterns of the battery storage. For all possible $2\cdot 3^{N}+2^{N}$ patterns, we perform the simulations from $t=kT_{\rm mini}$ to $t+T_{\rm pred}=(k+N)T_{\rm mini}$ based on the mathematical model of the distribution system given in Section II,
  • Step 2: Based on the simulation in Step 1, we compute the value of the evaluation function $J[k]$ for the possible $3^{N}$ switching patterns ${\bf\it\mathbf{\sigma}}_{\rm d}[k]$ at $t=kT_{\rm mini}$. We then compare the $3^{N}$ values of the functions. We choose the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}[k]$ that has the smallest value at $t=kT_{\rm mini}$. We explain the detailed procedure of the design of the optimal switching pattern in Subsection III-B.
  • Step 3: Based on the simulation and the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}^{*}[k]$ computed in Steps 1 and 2, respectively, we compute the evaluation function for the $2^{N}$ possible transmission patterns of the PV power generation equipment at $t=kT_{\rm mini}$. We compute the value of the evaluation function $J_{\rm pv}[k]$ for ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ of the transmission patterns of the PV power generation equipment at $t=kT_{\rm mini}$. Then, by comparing the $2^{N}$ values of the evaluation functions, we find the optimal transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ of the PV power generation equipment whose value is the minimum at $t=kT_{\rm mini}$. The detailed procedure of this step is explained in Subsection III-C.
  • Step 4: Based on the predictive simulation in Step 1, the optimal switching pattern ${\bf\it\mathbf{\sigma}}_{\rm d}^{*}[k]$ in Step 2, and the optimal transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,pv}[k]$ in Step 3, we compute the values of the evaluation function $J_{\rm battery}[k]$ for the $3^{N}$ battery storage transmission pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]$ at $t=kT_{\rm mini}$ for the evaluation function $J_{\rm battery}[k]$. In this case, the $3^{N}$ values of the evaluation function are compared, and the battery transfer pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}[k]$ with the minimum value at $t=kT_{\rm mini}$ is selected as the optimal power transfer pattern ${\bf\it\mathbf{\sigma}}_{\rm d,battery}^{*}[k]$. We explain the detailed procedure of the design of the optimal switching pattern in Subsection III-D.